OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and parallel to y = 2x + 4. y = -2x + 7 y = 2x - 3 y = 2x - 7 y = -2x + 3

OpenStudy (anonymous):

ok this is easy...What we are using is the point slope form okay?

OpenStudy (anonymous):

point slope form is (y-y0)=m(x-x0)...m=slope (x0.y0) the passing points

OpenStudy (anonymous):

to find the slope:parallel to y=2x+4 means our line has same slope as it

OpenStudy (anonymous):

looking at y=mx+c suggests that slope is 2

OpenStudy (anonymous):

so y=2x-7 is the equation of the line.....got it???@Trai7 ?

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

pleasure:)